OpenStudy (anonymous):

Solve x2 – 6x = 27 by completing the square

OpenStudy (anonymous):

\[x^2-6x=27\] \[(x-3)^2=27+3^2=27+9=36\] \[x-3=\pm\sqrt{36}=\pm6\] \[x-3=6\iff x=9\] \[x-3=-6\iff x=-3\]

OpenStudy (ash2326):

x^2-6x=27 x^2-6x-27=0 x^2-2*3x-27= adding and subtracting 9 x^2-2*3x+9-9-27=0 x^2-2*3x+3^2-36=0 (x-3)^2-6^2=0 a^2-b^2=(a-b)(a+b) (x-3+6)(x-3-6)=0 (x+3)(x-9)=0 x=-3 ,9